OpenStudy (anonymous):

Dillon has a bag of marbles. There are 3 blue, 1 green, 2 white, 4 black, and some red marbles in the bag. If the probability of pulling out a red marble is 3/4, how many red marbles are in the bag?

OpenStudy (anonymous):

3+1+2+4=10 so we have 1/4=10 so that means 1/4*4/1 = 4 3*10 = 30 so there are 30 red marbles in the bag... You see how I did that?

OpenStudy (anonymous):

or...x/(10+x)=(3/4) solve for x

OpenStudy (anonymous):

10=3+1+2+4